A person struggles to read by holding a book at arm’s length, a distance of 52 cm away. What power of reading glasses should be described for her if she wants to read at the “normal” near point of 25 cm? Assume the lens is very close to the eye.
1
Expert's answer
2020-05-05T18:42:49-0400

P = 1 / s = 1 / d0 + 1 / di = 1 / 0.25 - 1 / 0.55 = 2.18 D.